OpenStudy (anonymous):

what is The solution to 8x2 + 6x + 5 = 0

OpenStudy (anonymous):

Factoring?

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

what is 8x2?

OpenStudy (nincompoop):

is that 8 multiply by 2?

OpenStudy (anonymous):

^2 I Think

OpenStudy (anonymous):

no the variable x

OpenStudy (nincompoop):

so what is 2?

OpenStudy (anonymous):

the square root

OpenStudy (anonymous):

^2 Right?

OpenStudy (nincompoop):

\[ \sqrt{2} \]

OpenStudy (nincompoop):

that is a square root

OpenStudy (anonymous):

sorry i meant ^2

OpenStudy (nincompoop):

\[x^{\frac{m}{n}} = \sqrt[n]{x^m}\]

OpenStudy (nincompoop):

so what is ^2 what does that mean?

OpenStudy (anonymous):

as in 2^2 which would equal 4.... two squared

OpenStudy (nincompoop):

what have you studied about factoring thus far?

OpenStudy (anonymous):

its on like quadratic formulas and stuff

OpenStudy (nincompoop):

http://www.mathsisfun.com/algebra/factoring.html

OpenStudy (nincompoop):

then try what you've learned